OpenStudy (itsmichelle29):

help plz Convert the rectangular equation to a polar equation that expresses r in terms of thetaθ.

OpenStudy (itsmichelle29):

x^2 + (y-8)^2 = 64

zepdrix (zepdrix):

Ooo ok ok ok ok ok I got it.

zepdrix (zepdrix):

\[\large\rm x^2+(y-8)^2=64\]So recall that you use these two equations to make your conversion to polar,\[\large\rm x=r \cos \theta\qquad\qquad\qquad y=r \sin \theta\]So let's plug those in and see what happens,\[\large\rm (r \cos \theta)^2+(r \sin \theta-8)^2=64\]

zepdrix (zepdrix):

We replaced our `x with rcos(theta)` and `y with rsin(theta)`. Distribute the square in the first term. \[\large\rm r^2\cos^2\theta+(r \sin \theta-8)^2=64\]Do you understand how to expand out the other square?

OpenStudy (itsmichelle29):

nope not really